OpenStudy (anonymous):

what does 8-6.2 divided by 5 + 7 (0.91)

OpenStudy (dangerousjesse):

\(8-6.2=1.8\) \(7\times .91=6.37\) \(5+6.37=11.37\) \(1.8\div 11.37=?\)

OpenStudy (anonymous):

Thanks so much!!! Your a life saver! I hate math homework!